QUESTION 11

  1. When a firm uses cash to buy inventory, what happens to its current ratio?

    Current ratio decreases.

    Current ratio increases.

    Current ratio does not change.

    It cannot be determined.

1 points   

QUESTION 12

  1. Starbucks has cash of $2,757 million, inventory of $1,529 million, total current assets of $5,653 million, and total current liabilities of $6,168 million on its latest balance sheet. Compute its current ratio.

    0.45

    1.09

    0.92

    0.67

1 points   

QUESTION 13

  1. What is cash ratio?

    It is current assets / current liabilities.

    It is (current assets – inventory) / current liabilities.

    It is current liabilities / current assets.

    It is cash / current liabilities.

1 points   

QUESTION 15

  1. TJMaxx has the total assets of $14,326 million, total liabilities of $9,277 million, the earnings before interests and taxes of $4,218 million, interest expenses of $64 million, and depreciation of $817 million on its latest financial statements. Compute its equity multiplier.

    0.84

    1.84

    2.84

    1.12
Answer #1

11)Correct option is "C" - Current ratio does not change

when a firm buys inventory for cash ,Inventory is increased and cash is decreased .Since both are current asset ,so increase in one current asset (inventory) is offset by decrease in other current asset (cash) .Thus overall current ratio remains unchanged.

12)Correct option is "c"-.92

Current ratio =current asset /current liabilities

                  = 5653/6168

                   = .92

13)correct option is "D"

Cash ratio =[ Cash and cash equivalent + short term investment ]/current liabilities

15)correct option is "C"-2.84

Equity=Total asset -total liabilities

    = 14326-9277

       = 5049

Equity multiplier =Total asset /total equity

                      = 14326 / 5049

                      = 2.84